# Optimal convergence rates in trace distance and relative entropy for the quantum central limit theorem

### v1

@@ Abstract
A quantum analogue of the Central Limit Theorem (CLT), first introduced by Cushen and Hudson (1971), states that the $n$-fold convolution $\rho^{\boxplus n}$ of an $m$-mode quantum state $\rho$—with zero first moments and finite second moments—converges weakly, as $n$ increases, to a Gaussian state $\rho_{G}$ with the same first and second moments. Recently, this result has been extended with estimates of the convergence rate in various distance measures. In this paper, we establish optimal rates of convergence in both the trace distance and quantum relative entropy. Specifically, we show that for a centered $m$-mode quantum state with finite third-order moments, the trace distance between $\rho^{\boxplus n}$ and $\rho_{G}$ decays at the optimal rate of $\mathcal{O}(n^{-1/2})$, consistent with known convergence rates. Furthermore, for states with finite fourth-order moments (plus a small correction $\delta$ for $m>1$), we prove that the relative entropy between $\rho^{\boxplus n}$ and $\rho_{G}$ decays at the optimal rate of $\mathcal{O}(n^{-1})$. Both of these rates are proven to be optimal, even when assuming the finiteness of all moments of $\rho$. These results relax previous assumptions on higher-order moments, yielding convergence rates that match the best known results in the classical setting. Our proofs draw on techniques from the classical literature, including Edgeworth-type expansions of quantum characteristic functions, adapted to the quantum context. A key technical step in the proof of our entropic CLT is establishing an upper bound on the relative entropy distance between a general quantum state and its Gaussification. As a by-product of this, an upper bound on the relative entropy of non-Gaussianity is derived, which is of independent interest.
@@ 1 Introduction
As a landmark in probability theory, the Central Limit Theorem (CLT) states that the normalized sum $S_{n}=\frac{X_{1}+\cdots+X_{n}}{\sqrt{n}}$ of independent and identically distributed random variables $X_{1},X_{2},X_{3},\dots$ with zero mean converges to a Gaussian random variable $Z$ with the same mean and variance as that of $X_{1}$ [17]. Depending on the structure of the underlying distribution, this convergence can be characterized using various topologies or metrics. Two notable such CLTs are the Lindberg–Levy CLT [31], which asserts convergence in distribution under the assumption of a finite second moment of $X_{1}$, and the Berry–Esseen CLT [5], which asserts the uniform convergence of cumulative distribution functions at a rate of $\mathcal{O}(1/\sqrt{n})$, assuming the finiteness of the third absolute moment.
@@
In view of the wide range of applications of the Central Limit Theorem (CLT), several quantum (or non-commutative) analogues of the CLT have been proposed (see, e.g., [14, 25, 24, 19, 21, 33, 13, 26, 2, 11, 38, 35, 20, 27, 39, 22, 23, 12, 15, 30, 3]). In particular, Cushen and Hudson [14] introduced the first quantum CLT in the context of continuous-variable bosonic quantum systems. They demonstrated that if $\rho$ is a centered $m$-mode quantum state (i.e., its first-order moments vanish) with a finite covariance matrix, then the sequence of its $n$-fold symmetric convolutions $\rho^{\boxplus n}$ converges to a centered Gaussian state $\rho_{G}$ that has the same covariance matrix as $\rho$; this limiting state is referred to as the Gaussification of $\rho$. The definition of $\rho^{\boxplus n}$ will be provided in Section 2, but for now, we note that physically, the state $\rho^{\boxplus n}$ arises from the action of a sequence of beam splitters on $n$ identical copies of $\rho$, followed by tracing out all but the first $m$ output modes.
@@
The quantum CLT proposed by Cushen and Hudson was initially expressed in terms of convergence in the weak operator topology, which corresponds to pointwise convergence of the associated quantum characteristic functions. For quantum states, this topology coincides with the one induced by the trace distance (see [3, Section 3] and also [29, Lemma 4]). However, this theorem does not provide any bounds on the rate of convergence.
@@
In an effort to refine and strengthen the quantum CLT by establishing rates of convergence, Becker, Datta, Lami, and Rouzé [3] recently derived the asymptotic bound $\|\rho^{\boxplus n}-\rho_{G}\|_{2}=\mathcal{O}(1/\sqrt{n})$ for the convergence rate of the quantum CLT in terms of the Hilbert–Schmidt norm $\|\cdot\|_{2}$, assuming that the third moment of $\rho$ is finite. Building on this result, they also obtained the asymptotic bounds $\|\rho^{\boxplus n}-\rho_{G}\|_{1}=\mathcal{O}\big{(}n^{-1/2(m+1)}\big{)}$ and $D(\rho^{\boxplus n}\|\rho_{G})=\mathcal{O}\big{(}n^{-1/2(m+1)}\ln(n)\big{)}$ for the trace norm $\|\cdot\|_{1}$ and the quantum relative entropy $D(\cdot\|\cdot)$, under the same moment assumption.
@@
Subsequently, two authors of the current paper derived the asymptotic bounds $\|\rho^{\boxplus n}-\rho_{G}\|_{1}=\mathcal{O}(1/\sqrt{n})$ and $D(\rho^{\boxplus n}\|\rho_{G})=\mathcal{O}(1/n)$ in [4]. While these bounds align with the corresponding convergence rates in the classical setting and are shown to be optimal in [4], they are established under stronger assumptions. Specifically, the former bound is derived under the condition that the moment of order $\max\{3,2m\}$ is finite, whereas the latter relies on a Poincaré-type inequality in the quantum context, which is sufficiently strong to ensure the finiteness of all moments. When compared to classical results (with the exception of the single-mode case $m=1$), these assumptions appear quite stringent. Therefore, it is both important and highly desirable to relax these conditions.
@@
In this paper, we derive optimal rates of convergence for the quantum CLT with minimal assumptions. The first main result presented concerns the convergence rate in the trace distance.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite moment of order $3$. Let $\rho_{G}$ be the Gaussification of $\rho$. Then we have
@@
We note that this theorem for the single-mode case ($m=1$) has already been established in [4]; however, it is stronger than the corresponding result for the multi-mode case presented in that work. Furthermore, as illustrated by the example in [4], the convergence rate of $\mathcal{O}(1/\sqrt{n})$ mentioned above is tight, even under the assumption that all moments of $\rho$ are finite. Additionally, the assumptions made here align with the minimal moment conditions required to demonstrate the rate of $\mathcal{O}(1/\sqrt{n})$ in the classical context of the CLT regarding total variation distance.
The proof of Theorem 1 is based on the classical method of normal approximation, as detailed in [7], and incorporates ideas from [3, 4]. To estimate ${\big{\|}\rho^{\boxplus n}-\rho_{G}\big{\|}}_{1}$, we begin by projecting $\rho$ onto the subspace spanned by Fock states with a total particle number not exceeding $n$. This results in a truncated state of finite rank, where all moments remain finite. We then carry out a careful analysis of the moments and covariance matrix of this truncated state find controls on their asymptotic behavior as the threshold $n$ increases. Another key aspect of the proof involves extending the concept of Edgeworth-type expansions from classical literature to the quantum realm by analyzing an Edgeworth-type expansion for the quantum characteristic function. Ultimately, we leverage an estimate of the trace norm in terms of the Hilbert–Schmidt norm, along with the quantum Plancherel identity and Hölder’s inequality, to utilize the previously mentioned expansion and derive the desired bound.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite moment of order $4+\delta$, where $\delta=0$ if $m=1$, and $\delta>0$ otherwise. Let $\rho_{G}$ be the Gaussification of $\rho$. Then we have
@@
The proof of Theorem 2 draws inspiration from the approach developed by Bobkov, Chistyakov, and Götze in [10], where they link the relative entropy distance to a Sobolev norm of the characteristic functions. Their analysis involves examining higher-order derivatives of the characteristic function through Edgeworth-type expansions, yielding refined approximations of the norm. We follow a similar strategy. A key technical element in our proof is establishing an upper bound on the relative entropy distance between a general quantum state and its Gaussification, achieved at the operator level. We remark in passing that as an immediate corollary of this result, a general bound on the relative entropy of non-Gaussianity [18, 32] follows, which is of independent interest (see Section 5). Finally, by leveraging the quantum Plancherel identity and employing the truncation technique discussed earlier, we are able to perform a semiclassical analysis of the quantum characteristic function to obtain the desired bound.
The structure of this paper is as follows. In Section 2, we establish notation for bosonic quantum systems and phase space formalism and introduce the concept of quantum convolution in Subsection 2.3. In Section 3 we prove some basic results on moments of a quantum state and introduce the Edgeworth-type expansion for quantum states. Section 4 focuses on the truncation technique discussed earlier. In Section 5, we establish an upper bound on the relative entropy between a general quantum state and a Gaussian state in terms of the Hilbert–Schmidt norm. Sections 6 and 7 provide the proofs of Theorems 1 and 2, respectively, drawing on results from the preceding sections. Section 8 concludes with final remarks, and additional technical details are provided in the appendix.
@@ 2 Preliminaries
In this section, we review essential definitions and concepts regarding continuous-variable bosonic quantum systems, which will be used throughout the paper. Our primary reference for these topic is [36].
@@
The eigenstates of $N_{m}$ are the Fock states, denoted $|k_{1},\dots,k_{m}\rangle:=|k_{1}\rangle\otimes\cdots\otimes|k_{m}\rangle$, which form a complete orthonormal basis for $\mathcal{H}_{m}$. These Fock states, also known as number states, satisfy
@@ 2.2 Gaussian states
$\chi_{\sigma}(z)=\exp\Big{(}-\frac{1}{4}\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\sigma)\Lambda_{m}\hat{z}+i\mathbf{d}(\sigma)^{\top}\Lambda_{m}\hat{z}\Big{)},$ (14)
@@
A unitary $U$ is said to be Gaussian if it maps Gaussian states to Gaussian states. It is well-known that for any $m$-mode quantum state $\rho$, there exists a Gaussian unitary $U$ that both centers $\rho$ and diagonalizes its covariance matrix [36], i.e.
@@
The condition $\nu_{j}\geq 1$ follows from (13), since $\boldsymbol{\gamma}(\rho)$ is always positive definite with eigenvalues greater than or equal to 1. Notably, if $\nu_{j}=1$ (equivalently, $\beta_{j}=+\infty$), then $\tau_{j}=|0\rangle\langle 0|$ is the vacuum state. Specifically, $\nu_{j}=1$ implies that $\operatorname{tr}(\rho\mathbf{a}_{j}^{\dagger}\mathbf{a}_{j})=0$, so the support of the $j$-th mode of $\rho$ lies in the kernel of $\mathbf{a}_{j}$, which is spanned by the vacuum state. Thus, if $\rho$ is in the Williamson form and $\nu_{j}=1$, then the $j$-th mode is in the pure state $|0\rangle$ and is decoupled from other modes.
@@ 2.3 Quantum convolution
We define an addition rule for two $m$-mode quantum states using a specific Gaussian unitary associated with a beam splitter [28]. Let $\rho$ and $\sigma$ be two $m$-mode quantum states, and let $0\leq\eta\leq 1$. The quantum convolution of $\rho$ and $\sigma$ with parameter $\eta$ is defined as
@@
$U_{\eta}:=\exp\left(\arccos(\sqrt{\eta})\sum_{j=1}^{m}\big{(}\mathbf{a}_{j,1}^{\dagger}\mathbf{a}_{j,2}-\mathbf{a}_{j,1}\mathbf{a}_{j,2}^{\dagger}\big{)}\right).$
Here, $U_{\eta}$ is a unitary operator acting on $2m$ modes, which we treat as two $m$-mode subsystems. The operators $\mathbf{a}_{j,1}$ and $\mathbf{a}_{j,2}$ for $j=1,\dots,m$ denote the annihilation operators on the first and second subsystems, respectively, and $\mathrm{tr}_{2}(\cdot)$ represents the partial trace over the second subsystem.
The Gaussian unitary operator $U_{\eta}$ is equivalently described by its action on the annihilation and creation operators:
@@
where $\delta$ depends on $\epsilon$ and the second moment of $\rho$ (see [3, Proposition 15]). Interestingly, this property does not hold in such a general form for classical probability distributions. Using this property along with (21), we see that for $\|z\|>\epsilon\sqrt{n}$, $|\chi_{\rho^{\boxplus n}}(z)|\leq(1-\delta)^{n}$ becomes exponentially small. This inequality is critical in our analysis of convergence rates in the quantum CLT.

### v2

@@ Abstract
A quantum analogue of the Central Limit Theorem (CLT), first introduced by Cushen and Hudson (1971), states that the $n$-fold convolution $\rho^{\boxplus n}$ of an $m$-mode quantum state $\rho$—with zero first moments and finite second moments—converges weakly, as $n$ increases, to a Gaussian state $\rho_{G}$ with the same first and second moments. Recently, this result has been extended with estimates of the convergence rate in various distance measures. In this paper, we establish optimal rates of convergence in both the trace distance and quantum relative entropy. Specifically, we show that for a centered $m$-mode quantum state with finite third-order moments, the trace distance between $\rho^{\boxplus n}$ and $\rho_{G}$ decays at the optimal rate of $\mathcal{O}(n^{-1/2})$. Furthermore, for states with finite fourth-order moments (order $4+\delta$ for an arbitrary small $\delta>0$ if $m>1$), we prove that the relative entropy between $\rho^{\boxplus n}$ and $\rho_{G}$ decays at the optimal rate of $\mathcal{O}(n^{-1})$. Both of these rates are proven to be optimal, even when assuming the finiteness of all moments of $\rho$. These results relax previous assumptions on higher-order moments, yielding convergence rates that match the best known results in the classical setting. By giving explicit examples we also show that our moment assumptions are essentially minimal. We show that for any $\theta>0$, there exists a quantum state $\rho$ with finite moments of order less than $3-\theta$, such that the convergence rate of $\rho^{\boxplus n}$ to $\rho_{G}$ in trace distance is not $\mathcal{O}(n^{-1/2})$. Similarly, we show that for any $\theta>0$, there exists a quantum state $\rho$ with finite moments of order less than $4-\theta$, such that the relative entropy between $\rho^{\boxplus n}$ to $\rho_{G}$ does not decay at the rate $\mathcal{O}(n^{-1})$. Our proofs draw on techniques from the classical literature, including Edgeworth-type expansions of quantum characteristic functions, adapted to the quantum context. A key technical step in the proof of our entropic CLT is establishing an upper bound on the relative entropy distance between a general quantum state and its Gaussification, which is of independent interest.
@@ 1 Introduction
As a landmark in probability theory, the Central Limit Theorem (CLT) states that the normalized sum $S_{n}=\frac{X_{1}+\cdots+X_{n}}{\sqrt{n}}$ of independent and identically distributed random variables $X_{1},X_{2},X_{3},\dots$ with zero mean, converges to a Gaussian random variable $Z$ with the same mean and variance as that of $X_{1}$ [16]. Depending on the structure of the underlying distribution, this convergence can be characterized using various topologies or metrics. Two notable such CLTs are the Lindberg–Levy CLT [31], which asserts convergence in distribution under the assumption of a finite second moment of $X_{1}$, and the Berry–Esseen CLT [5], which asserts the uniform convergence of cumulative distribution functions at a rate of $\mathcal{O}(1/\sqrt{n})$, assuming the finiteness of the third absolute moment.
@@
In view of the wide range of applications of the CLT, several quantum (or non-commutative) analogues of the CLT have been proposed (see, e.g., [13, 24, 23, 18, 20, 33, 12, 26, 2, 38, 35, 19, 27, 39, 21, 22, 11, 14, 30, 3]). In particular, Cushen and Hudson [13] introduced the first quantum CLT in the context of continuous-variable bosonic quantum systems. They demonstrated that if $\rho$ is a centered $m$-mode quantum state (i.e., its first-order moments vanish) with a finite covariance matrix, then the sequence of its $n$-fold symmetric convolutions $\rho^{\boxplus n}$ converges to a centered Gaussian state $\rho_{G}$ that has the same covariance matrix as $\rho$; this limiting state is referred to as the Gaussification of $\rho$. The definition of $\rho^{\boxplus n}$ will be provided in Section 2, but for now, we note that physically, the state $\rho^{\boxplus n}$ arises from the action of a sequence of beam splitters on $n$ identical copies of $\rho$, followed by tracing out all but the first $m$ output modes.
@@
The quantum CLT proposed by Cushen and Hudson was initially expressed in terms of convergence in the weak operator topology, which corresponds to pointwise convergence of the associated quantum characteristic functions. For quantum states, this topology coincides with the one induced by the trace distance (see [3, Section 3] and  [29, Lemma 4]). However, this theorem does not provide any bounds on the rate of convergence.
@@
In an effort to refine and strengthen the quantum CLT by establishing rates of convergence, Becker, Datta, Lami, and Rouzé [3] recently derived the asymptotic bound $\|\rho^{\boxplus n}-\rho_{G}\|_{2}=\mathcal{O}(1/\sqrt{n})$ for the convergence rate of the quantum CLT in terms of the Hilbert–Schmidt norm $\|\cdot\|_{2}$, assuming that the third-order moments of $\rho$ are finite. Building on this result, they also obtained the asymptotic bounds $\|\rho^{\boxplus n}-\rho_{G}\|_{1}=\mathcal{O}\big{(}n^{-1/2(m+1)}\big{)}$ and $D(\rho^{\boxplus n}\|\rho_{G})=\mathcal{O}\big{(}n^{-1/2(m+1)}\ln(n)\big{)}$ for the trace norm $\|\cdot\|_{1}$ and the quantum relative entropy $D(\cdot\|\cdot)$, under the same moment assumption.
@@
Subsequently, two authors of the current paper derived the asymptotic bounds $\|\rho^{\boxplus n}-\rho_{G}\|_{1}=\mathcal{O}(1/\sqrt{n})$ and $D(\rho^{\boxplus n}\|\rho_{G})=\mathcal{O}(1/n)$ in [4]. While these bounds align with the corresponding convergence rates in the classical setting and are shown to be optimal in [4], they are established under stronger assumptions. Specifically, the former bound is derived under the condition that the moments of order $\max\{3,2m\}$ is finite, whereas the latter relies on a Poincaré-type inequality in the quantum context, which is sufficiently strong to ensure the finiteness of all moments. When compared to classical results (with the exception of the single-mode case $m=1$), these assumptions appear quite stringent. Therefore, it is both important and highly desirable to relax these conditions.
@@
In this paper, we derive optimal rates of convergence for the quantum CLT with minimal assumptions. Our first main result concerns the convergence rate in the trace distance.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite moments of order $3$. Let $\rho_{G}$ be the Gaussification of $\rho$. Then we have
@@
We note that this theorem for the single-mode case ($m=1$) has already been established in [4]; however, it is stronger than the corresponding result for the multi-mode case presented in that work. Furthermore, as illustrated by the example in [4], the convergence rate of $\mathcal{O}(1/\sqrt{n})$ mentioned above is tight, even under the assumption that all moments of $\rho$ are finite. Additionally, the assumptions made here aligned with the minimal moment conditions required to demonstrate the convergence rate of $\mathcal{O}(1/\sqrt{n})$ in the classical CLT regarding total variation distance.
The proof of Theorem 1 is based on the classical method of normal approximation, as detailed in [7], and incorporates ideas from [3, 4]. To estimate ${\big{\|}\rho^{\boxplus n}-\rho_{G}\big{\|}}_{1}$, we begin by projecting $\rho$ onto the subspace spanned by Fock states with a total particle number not exceeding $n$. This results in a truncated state of finite rank, where all moments remain finite. We then carry out a careful analysis of the moments and covariance matrix of this truncated state and find controls on their asymptotic behavior as $n$ increases. Another key aspect of the proof involves extending the concept of Edgeworth-type expansions from the classical literature to the quantum realm by analyzing an Edgeworth-type expansion for the quantum characteristic function. Ultimately, we leverage an estimate of the trace norm in terms of the Hilbert–Schmidt norm, along with the quantum Plancherel identity and Hölder’s inequality, to utilize the previously mentioned expansion and derive the desired bound.
@@
Let $\rho$ be a centered $m$-mode quantum state with finite moments of order $4+\delta$, where $\delta=0$ if $m=1$, and $\delta>0$ otherwise. Let $\rho_{G}$ be the Gaussification of $\rho$. Then we have
@@
The proof of Theorem 2 draws inspiration from the approach developed by Bobkov, Chistyakov, and Götze in [10], where they link the relative entropy distance to a Sobolev norm of the characteristic functions. Their analysis involves examining higher-order derivatives of the characteristic function through Edgeworth-type expansions, yielding refined approximations of the norm. We follow a similar strategy. A key technical element in our proof is establishing an upper bound on the relative entropy distance between a general quantum state and its Gaussification, achieved at the operator level. We remark in passing that as an immediate corollary of this result, a general bound on the relative entropy of non-Gaussianity [17, 32] follows, which is of independent interest. Finally, by leveraging the quantum Plancherel identity and employing the truncation technique discussed earlier, we are able to perform a semiclassical analysis of the quantum characteristic function to obtain the desired bound.
We also by giving explicit examples show that the moment assumptions in Theorems 1 and 2 are essentially minimal and cannot be weakened.
Theorem 3.
For any $10<\theta<$ 1 the followings hold:
  1. (i)
    There is a single-mode quantum state $\rho$ whose moments of order $\kappa$, for any $\kappa<3-\theta$, are finite, while
    $\sup_{n}\sqrt{n}\big{\|}\rho^{\boxplus n}-\rho_{G}\big{\|}_{1}=+\infty.$
  2. (ii)
    There is a single-mode quantum state $\rho$ whose moments of order $\kappa$, for any $\kappa<4-\theta$, are finite, while
    $\sup_{n}nD\big{(}\rho^{\boxplus n}\big{\|}\rho_{G}\big{)}=+\infty.$
To construct these examples, we employ ideas from the classical literature, particularly techniques from [25] for the trace distance, and from [9] for the relative entropy.
The structure of this paper is as follows. In Section 2, we establish notation for bosonic quantum systems and phase space formalism and introduce the concept of quantum convolution in Subsection 2.3. In Section 3 we prove some basic results on moments of a quantum state and introduce the Edgeworth-type expansion for quantum states. Section 4 focuses on the truncation technique discussed earlier. In Section 5, we establish an upper bound on the relative entropy between a general quantum state and a Gaussian state in terms of the Hilbert–Schmidt norm. Sections 6 and 7 provide the proofs of Theorems 1 and 2, respectively, drawing on results from the preceding sections. The examples showing the essential minimality of our assumptions on the moments are presented in Section 8. Section 9 concludes with final remarks, and additional technical details are provided in the appendices.
@@ 2 Preliminaries
In this section, we review essential definitions and concepts regarding continuous-variable bosonic quantum systems, which will be used throughout the paper. Our primary reference for this topic is [36].
@@
The eigenstates of $N_{m}$ are the Fock states, denoted by $|k_{1},\dots,k_{m}\rangle:=|k_{1}\rangle\otimes\cdots\otimes|k_{m}\rangle$, which form a complete orthonormal basis for $\mathcal{H}_{m}$. These Fock states, also known as number states, satisfy
@@ 2.1 Phase space formalism
Obviously, we can write the characteristic function using Wigner function as
$\chi_{T}(z)=\int W_{T}(u)e^{u^{\dagger}z-z^{\dagger}u}\text{\rm{d}}^{2m}u$ (11)
@@ 2.2 Gaussian states
$\chi_{\sigma}(z)=\exp\bigg{(}-\frac{1}{4}\hat{z}^{\dagger}\Lambda_{m}^{\dagger}\boldsymbol{\gamma}(\sigma)\Lambda_{m}\hat{z}+i\mathbf{d}(\sigma)^{\top}\Lambda_{m}\hat{z}\bigg{)},$ (15)
@@
A unitary $U$ is said to be Gaussian if it maps Gaussian states to Gaussian states under conjugation. It is well-known that for any $m$-mode quantum state $\rho$, there exists a Gaussian unitary $U$ that both centers $\rho$ and diagonalizes its covariance matrix [36], i.e.
@@
The condition $\nu_{j}\geq 1$ follows from (14), since $\boldsymbol{\gamma}(\rho)$ is always positive definite with eigenvalues greater than or equal to 1. Notably, if $\nu_{j}=1$ (equivalently, $\beta_{j}=+\infty$), then $\tau_{j}=|0\rangle\langle 0|$ is the vacuum state. Specifically, $\nu_{j}=1$ implies that $\operatorname{tr}(\rho\mathbf{a}_{j}^{\dagger}\mathbf{a}_{j})=0$, so the support of the $j$-th mode of $\rho$ lies in the kernel of $\mathbf{a}_{j}$, which is spanned by the vacuum state. Thus, if $\rho$ is in the Williamson form and $\nu_{j}=1$, then the $j$-th mode is in the pure state $|0\rangle$ and is decoupled from the other modes.
@@ 2.3 Quantum convolution
We define an addition rule for two $m$-mode quantum states using a specific Gaussian unitary associated with a beam splitter (see [28] and references therein). Let $\rho$ and $\sigma$ be two $m$-mode quantum states, and let $0\leq\eta\leq 1$. The quantum convolution of $\rho$ and $\sigma$ with parameter $\eta$ is defined as
@@
$U_{\eta}:=\exp\bigg{(}\arccos(\sqrt{\eta})\sum_{j=1}^{m}\big{(}\mathbf{a}_{j,1}^{\dagger}\mathbf{a}_{j,2}-\mathbf{a}_{j,1}\mathbf{a}_{j,2}^{\dagger}\big{)}\bigg{)}.$
Here, $U_{\eta}$ is a unitary operator acting on $2m$ modes, which we treat as two $m$-mode subsystems. The operators $\mathbf{a}_{j,1}$ and $\mathbf{a}_{j,2}$ for $j=1,\dots,m$ denote the annihilation operators on the first and second subsystems, respectively, and $\mathrm{tr}_{2}(\cdot)$ represents the partial trace over the second subsystem, consisting of $m$ modes. The Gaussian unitary operator $U_{\eta}$ is equivalently described by its action on the annihilation and creation operators:
@@
where $\delta$ depends on $\epsilon$ and the covariance matrix of $\rho$ (see [3, Proposition 15]). Interestingly, this property does not hold in such a general form for classical probability distributions. Using this property along with (22), we see that for $\|z\|>\epsilon\sqrt{n}$, $|\chi_{\rho^{\boxplus n}}(z)|\leq(1-\delta)^{n}$ becomes exponentially small. This inequality is critical in our analysis of convergence rates in the quantum CLT.